# Simulating Polynomial-Time Nondeterministic Turing Machines via Nondeterministic Turing Machines

### v21

@@ Abstract.
We prove in this paper that there is a language $L_{s}$ accepted by some nondeterministic Turing machine that runs within time $O(n^{k})$ for any positive integer $k\in\mathbb{N}_{1}$ but not by any ${\rm co}\mathcal{NP}$ machines. Then we further show that $L_{s}$ is in $\mathcal{NP}$, thus proving that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$. The main techniques used in this paper are simulation and the novel new techniques developed in the author’s recent work. Our main result has profound implications, such as $\mathcal{P}\neq\mathcal{NP}$, etc. Further, if there exists some oracle $A$ such that $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, we then explore what mystery lies behind it and show that if $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, then the set of all ${\rm co}\mathcal{NP}^{A}$ machines is not enumerable, thus showing that the simulation techniques are not applicable for the first half of the whole step to separate $\mathcal{NP}^{A}$ from ${\rm co}\mathcal{NP}^{A}$. Finally, a lower bounds result for Frege proof systems is presented (i.e., no Frege proof systems can be polynomially bounded).
@@ 1. Introduction
Our motivation in this paper is that, in the author’s recent work [Lin21], where we presented a proof that $\mathcal{P}\neq\mathcal{NP}$, we were completely at a loss and didn’t know how to prove $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, so we left this outstanding and long-standing conjecture of $\mathcal{NP}\neq{\rm co}\mathcal{NP}$ untouched. In the present paper, we try to resolve this important open conjecture.
@@ 1.1. Our Contributions
In this paper, we explore and settle the aforementioned open problems. Our first main goal in this paper is to prove the following most important theorem:
@@
As mentioned earlier, Theorem 1.1 has profound implications, one of which is the following important corollary.  [^1] In addition, if the complexity class $\mathcal{BQP}$ (bounded-error quantum polynomial-time; see page 213 of [AB09]) has the similar property (i.e., $\mathcal{BQP}=\rm{co}\mathcal{BQP}$), then obviously we have another important relation $\mathcal{BQP}\neq\mathcal{NP}$.
@@
Since $\mathcal{EXP}={\rm co}\mathcal{EXP}$ (where $\mathcal{EXP}$ is the class of languages accepted by deterministic Turing machines in time $2^{n^{O(1)}}$; see Section 2 for big $O$ notation), this, together with Corollary 1.4, follows that $\mathcal{EXP}\neq\mathcal{NEXP}$. Furthermore, the complexity class $\mathcal{BPP}$ (bounded-error probabilistic polynomial-time) aims to capture the set of decision problems efficiently solvable by polynomial-time probabilistic Turing machines (see e.g., [AB09]), and a central open problem in probabilistic complexity theory is the relation between $\mathcal{BPP}$ and $\mathcal{NEXP}$. However, currently researchers only know that $\mathcal{BPP}$ is sandwiched between $\mathcal{P}$ and $\mathcal{EXP}$ (i.e., $\mathcal{P}\subseteq\mathcal{BPP}\subseteq\mathcal{EXP}$), but are even unable to show that $\mathcal{BPP}$ is a proper subset of $\mathcal{NEXP}$; see e.g., page 126 of [AB09]. With the consequence of Corollary 1.4 (i.e., $\mathcal{EXP}\neq\mathcal{NEXP}$) at hand, it immediately follows that
@@
It is interesting that the complexity class $\mathcal{NP}$ has a rich structure if $\mathcal{P}$ and $\mathcal{NP}$ differ. Specifically, in [Lad75], Lander constructed a language that is $\mathcal{NP}$-intermediate under the assumption that $\mathcal{P}\neq\mathcal{NP}$. In fact, since by Theorem 1.1, we know that $\mathcal{P}\neq{\rm co}\mathcal{NP}$ and symmetrically also have the following interesting outcome saying that the complexity class ${\rm co}\mathcal{NP}$ has the intermediate languages. To see so, let $L\in\mathcal{NP}$ be the $\mathcal{NP}$-intermediate language constructed in [Lad75]; then $\overline{L}=\{0,1\}^{*}-L$ is a ${\rm co}\mathcal{NP}$-intermediate language in ${\rm co}\mathcal{NP}$. In other words, we will prove in detail the following important result:
@@ 1.2. Background and Prior Work
In history, the fundamental measure of time opened the door to the study of the extremely expressive time complexity class $\mathcal{NP}$, one of the most important classical complexity classes, i.e., nondeterministic polynomial-time. Specifically, $\mathcal{NP}$ is the set of decision problems for which the problem instances, where the answer is “yes,” have proofs verifiable in polynomial time by some deterministic Turing machine, or alternatively the set of problems that can be solved in polynomial time by a nondeterministic Turing machine (see e.g., [A2]). The famous Cook-Levin theorem [Coo71, Lev73] shows that this class has complete problems, which states that the Satisfiability is $\mathcal{NP}$-complete, i.e., Satisfiability is in $\mathcal{NP}$ and any other language in $\mathcal{NP}$ can be reduced to it in polynomial time. At the same time, it is also worth noting that this famous and seminal result opened the door to research into the renowned rich theory of $\mathcal{NP}$-completeness [Kar72].
@@
In the next few paragraphs, let us recall some of the interesting background, the current status, the history, and the main goals of the field of proof complexity. On the one hand, as we all know well, proof theory is a major branch of mathematical logic and theoretical computer science with which proofs are treated as formal mathematical objects, facilitating their analysis by mathematical techniques (see e.g., [A4]). Some of the major areas of proof theory include structural proof theory, ordinal analysis, automated theorem proving, and proof complexity, and so forth. It should be pointed out that much research also focuses on applications in computer science, linguistics, and philosophy, from which we see that the impact of proof theory is very profound. Moreover, as pointed out in [Raz04] by Razborov, proof and computations are among the most fundamental concepts relevant to virtually any human intellectual activity. Both have been central to the development of mathematics for thousands of years. The effort to study these concepts themselves in a rigorous, metamathematical way was initiated in the $20$th century and led to the flourishing of mathematical logic and derived disciplines.
@@
Likewise, propositional proof complexity is an area of study that has seen a rapid development over the last two decades, which plays as important a role in the theory of feasible proofs as the role played by the complexity of Boolean circuits in the theory of efficient computations; see e.g., the celebrated work [Raz15] by Razborov. In most cases, according to [Raz15], the basic question of propositional proof complexity can be reduced to that given a mathematical statement encoded as a propositional tautology $\psi$ and a class of admissible mathematical proofs formalized as a propositional proof system $P$, what is the minimal possible complexity of a $P$-proof of $\psi$? In other words, propositional proof complexity aims to understand and analyze the computational resources required to prove propositional tautologies, in the same way that circuit complexity studies the resources required to compute Boolean functions. In light of the point of view given in [Raz15], the task of proving lower bounds for strong proof systems like Frege or Extended Frege modulo any hardness assumption in the purely computational world may be almost as interesting; see e.g., [Raz15]. It is worth noting that it has been observed by Razborov [Raz15] that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$ implies lower bounds for any propositional proof system. In addition, see e.g., [Raz03] for interesting directions in propositional proof complexity.
@@
Let us return back to the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem for this moment. As a matter of fact, the so-called $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem is the central problem in proof complexity; see e.g., the excellent survey [Kra19]. It formalizes the question of whether or not there are efficient ways to prove the negative cases in our and in many other similar examples. Moreover, the ultimate goal of proof complexity is to show that there is no universal propositional proof system allowing for efficient proofs of all tautologies, which is equivalent to showing that the computational complexity class $\mathcal{NP}$ is not closed under the complement; see e.g., [Kra19].
Particularly, in 1974, Cook and Reckhow showed in [CR74] that there exists a super proof system if and only if $\mathcal{NP}$ is closed under complement; that is, if and only if $\mathcal{NP}={\rm co}\mathcal{NP}$. It is, of course, open whether any superproof system exists. The above result has led to what is sometimes called “Cook’s program” for proving $\mathcal{NP}\neq{\rm co}\mathcal{NP}$: prove superpolynomial lower bounds for proof lengths in stronger and stronger proposition proof systems until they are established for all abstract proof systems. Cook’s program, which also can be seen as a program to prove $\mathcal{P}\neq\mathcal{NP}$, is an attractive and plausible approach; unfortunately, it has turned out to be quite hard to establish superpolynomial lower bounds on common proof systems; see e.g., the excellent survey on proof complexity [Bus12].
@@ 1.4. Organization
The remainder of this paper is organized as follows: For the convenience of the reader, we will review some notions in the next section closely associated with our discussions and fix some notation we will use in the following context, with some useful technical lemmas also presented in the next section. In Section 3, we prove that all polynomial-time nondeterministic Turing machines are enumerable and all ${\rm co}\mathcal{NP}$-machines are enumerable, thus showing that there exists an enumeration of all ${\rm co}\mathcal{NP}$-machines. Section 4 contains the definition of our universal nondeterministic Turing machine, which accepts a language $L_{s}$ not in ${\rm co}\mathcal{NP}$, showing the desired lower bounds. The Section 5 is devoted to showing that the language $L_{s}$ is in fact in $\mathcal{NP}$, proving the upper bounds required, in which we finish the proof of Theorem 1.1 at our own convenience. In Section 6 we will show Theorem 1.3 to break the so-called Relativization Barrier. In Section 7, we show that there are ${\rm co}\mathcal{NP}$-intermediate languages in the complexity class ${\rm co}\mathcal{NP}$. In Section 8, we prove Theorem 1.8, which says that no Frege proof systems are polynomially bounded, thus answering an important open question in the area of proof complexity. Finally, we list some open problems for future research in the last section.
@@ 2. Preliminaries
Let $\mathbb{N}_{1}=\{1,2,3,\cdots\}$, where $\infty\notin\mathbb{N}_{1}$, i.e., the set of all positive integers. We also denote by $\mathbb{N}$ the set of natural numbers,333 For the natural number system and $\infty$ (i.e., $\infty$ here), we refer the reader to the interesting textbook [Tao22], in which Tao provides a full and philosophical treatment of them. i.e.,
@@
means that there exist a positive integer $N_{0}$ and a positive constant $M$ such that
@@
The little $o$ notation also indicates the order of growth of some quantity as a function of $n$ or the limiting behavior of a function, but with different meaning. Specifically, that $T(n)$ is little $o$ of $t(n)$, i.e.,
@@
Throughout this paper, the computational modes used are nondeterministic Turing machines (or their variants such as nondeterministic Turing machines with oracle). We follow the standard definition of a nondeterministic Turing machine given in the standard textbook [AHU74]. Let us first introduce the precise definition of a nondeterministic Turing machine as follows:
@@
Let $M$ be a nondeterministic Turing machine, and $w$ be an input. Then $M(w)$ represents that $M$ is on input $w$.
@@
Let $M$ be a polynomial-time (say, $t(n)$ time-bounded) nondeterministic Turing machine over the input alphabet $\{0,1\}$. Then, under the “there exists no accepting path” condition, a word $w$ is rejected by $M$ if there exists no computation path $\gamma$ putting the machine into an accepting state on input $w$ (in such a situation, we call $M$ rejects $w$). Clearly, $M$ will define the language $\overline{L}(M)$, which is the complement of $L(M)$ (i.e., $\overline{L}(M)=\{0,1\}^{*}\setminus L(M)$). Specifically, $w\in\overline{L}(M)$ if and only if $M$ rejects $w$.
Since ${\rm co}\mathcal{NP}=\{\overline{L}\,:\,L\in\mathcal{NP}\}$, a polynomial-time nondeterministic Turing machine $M$ accepts the language $L$ under the condition of “there exists at least one accepting path,” the same machine $M$ will simultaneously define the language $\overline{L}$ under the condition of “there exists no accepting path.” Thus, we can conveniently define the notion of a ${\rm co}\mathcal{NP}$ machine as follows:
@@
A ${\rm co}\mathcal{NP}$ machine $M$ is exactly a polynomial-time nondeterministic Turing machine $M$ working under the “there exists no accepting path” condition. We define the language accepted by the ${\rm co}\mathcal{NP}$ machine $M$ to be $\overline{L}(M)$. 444 In short, there is an interesting relationship: ${\rm co}\mathcal{NP}$ machine $M$ accepts the language $\overline{L}(M)$, but polynomial-time nondeterministic Turing machine $M$ rejects $\overline{L}(M)$; polynomial-time nondeterministic Turing machine $M$ accepts the language $L(M)$, but ${\rm co}\mathcal{NP}$ machine $M$ rejects $L(M)$. Namely, for any input $w$, $w\in\overline{L}(M)$ if and only if polynomial-time nondeterministic Turing machine $M$ rejects $w$.
Specifically, assume that the $t(n)$ time-bounded ${\rm co}\mathcal{NP}$ machine $M$ (over the input alphabet $\{0,1\}$) accepts the language $\overline{L}(M)$. Clearly, $w\notin\overline{L}(M)$ if and only if polynomial-time nondeterministic Turing machine $M$ accepts $w$.

### v22

@@ Abstract.
We prove in this paper that there is a language $L_{s}$ accepted by some nondeterministic Turing machine that runs within time $O(n^{k})$ for any positive integer $k\in\mathbb{N}_{1}$ but not by any ${\rm co}\mathcal{NP}$ machines. Then we further show that $L_{s}$ is in $\mathcal{NP}$, thus proving that
$\mathcal{NP}\neq{\rm co}\mathcal{NP}.$
The main techniques used in this paper are simulation and the novel new techniques developed in the author’s recent work. Our main result has profound implications, such as $\mathcal{P}\neq\mathcal{NP}$, etc. Further, if there exists some oracle $A$ such that $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, we then explore what mystery lies behind it and show that if $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, then the set of all ${\rm co}\mathcal{NP}^{A}$ machines is not enumerable, thus showing that the simulation techniques are not applicable for the first half of the whole step to separate $\mathcal{NP}^{A}$ from ${\rm co}\mathcal{NP}^{A}$. Finally, a lower bounds result for Frege proof systems is presented (i.e., no Frege proof systems can be polynomially bounded).
@@ 1. Introduction
Our motivation in this paper is that, in the author’s recent work [Lin21], where we presented a proof that $\mathcal{P}\neq\mathcal{NP}$, we were completely at a loss and didn’t know how to prove $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, so we left this outstanding and long-standing conjecture of $\mathcal{NP}\neq{\rm co}\mathcal{NP}$ untouched. In the present paper, we try to resolve this important open conjecture. We would like to stress that, generally, the conjecture of $\mathcal{NP}\neq{\rm co}\mathcal{NP}$ is believed to be harder than the conjecture of $\mathcal{P}\neq\mathcal{NP}$, since the former easily implies the latter.
@@ 1.1. Our Contributions
In this paper, we explore and settle the aforementioned important open problems. Our first main goal in this paper is to prove the following most important theorem:
@@
As mentioned earlier, Theorem 1.1 has profound implications, one of which is the following important corollary.
@@
Moreover, the complexity class $\mathcal{BQP}$ (bounded-error quantum polynomial-time; see page 213 of [AB09]) has the similar property like $\mathcal{P}={\rm co}\mathcal{P}$ (i.e., $\mathcal{BQP}=\rm{co}\mathcal{BQP}$), then obviously we have the first important conclusion regarding the important relations between the complexity classes $\mathcal{BQP}$ and $\mathcal{NP}$:
Corollary 1.4.
$\mathcal{BQP}\neq\mathcal{NP}$. [^2] But the precise relationships between $\mathcal{BQP}$ and $\mathcal{NP}$ are still open; see e.g., Section 9.
@@
Since $\mathcal{EXP}={\rm co}\mathcal{EXP}$ (where $\mathcal{EXP}$ is the class of languages accepted by deterministic Turing machines in time $2^{n^{O(1)}}$; see Section 2 for big $O$ notation), this, together with Corollary 1.5, implies that $\mathcal{EXP}\neq\mathcal{NEXP}$. Furthermore, the complexity class $\mathcal{BPP}$ (bounded-error probabilistic polynomial-time) aims to capture the set of decision problems efficiently solvable by polynomial-time probabilistic Turing machines (see e.g., [AB09]), and a central open problem in probabilistic complexity theory is the relation between $\mathcal{BPP}$ and $\mathcal{NEXP}$. However, currently researchers only know that $\mathcal{BPP}$ is sandwiched between $\mathcal{P}$ and $\mathcal{EXP}$ (i.e., $\mathcal{P}\subseteq\mathcal{BPP}\subseteq\mathcal{EXP}$), but are even unable to show that $\mathcal{BPP}$ is a proper subset of $\mathcal{NEXP}$; see e.g., page 126 of [AB09]. With the consequence of Corollary 1.5 (i.e., $\mathcal{EXP}\neq\mathcal{NEXP}$) at hand, it immediately follows that
@@
It is interesting that the complexity class $\mathcal{NP}$ has a rich structure if $\mathcal{P}$ and $\mathcal{NP}$ differ. Specifically, in [Lad75], Lander constructed a language that is $\mathcal{NP}$-intermediate under the assumption that $\mathcal{P}\neq\mathcal{NP}$. In fact, since by Theorem 1.1, we know that $\mathcal{P}\neq{\rm co}\mathcal{NP}$, and symmetrically, we also have the following interesting outcome saying that the complexity class ${\rm co}\mathcal{NP}$ has the intermediate languages. To see so, let $L\in\mathcal{NP}$ be the $\mathcal{NP}$-intermediate language constructed in [Lad75]; then $\overline{L}=\{0,1\}^{*}\setminus L$ is a ${\rm co}\mathcal{NP}$-intermediate language in ${\rm co}\mathcal{NP}$. In other words, we will prove in detail the following important result:
@@ 1.2. Background and Prior Work
In history, the fundamental measure of time opened the door to the study of the extremely expressive time complexity class $\mathcal{NP}$, one of the most important classical complexity classes, i.e., nondeterministic polynomial-time. Specifically, $\mathcal{NP}$ is the set of decision problems for which the problem instances, where the answer is “yes,” have proofs verifiable in polynomial time by some deterministic Turing machine, or alternatively the set of problems that can be solved in polynomial time by a nondeterministic Turing machine (see e.g., [A2]). The famous Cook-Levin theorem [Coo71, Lev73] shows that this class has complete problems, which states that Satisfiability is $\mathcal{NP}$-complete, i.e., Satisfiability is in $\mathcal{NP}$ and any other language in $\mathcal{NP}$ can be reduced to it in polynomial time. At the same time, it is also worth noting that this famous and seminal result opened the door to research into the renowned rich theory of $\mathcal{NP}$-completeness [Kar72].
@@
In the next few paragraphs, let us recall some of the interesting background, the current status, the history, and the main goals of the field of proof complexity. On the one hand, as we all know well, proof theory is a major branch of mathematical logic and theoretical computer science in which proofs are treated as formal mathematical objects, facilitating their analysis by mathematical techniques (see e.g., [A4]). Some of the major areas of proof theory include structural proof theory, ordinal analysis, automated theorem proving, and proof complexity, and so forth. It should be pointed out that much research also focuses on applications in computer science, linguistics, and philosophy, from which we see that the impact of proof theory is very profound. Moreover, as pointed out in [Raz04] by Razborov, proof and computations are among the most fundamental concepts relevant to virtually any human intellectual activity. Both have been central to the development of mathematics for thousands of years. The effort to study these concepts themselves in a rigorous, metamathematical way was initiated in the $20$th century and led to the flourishing of mathematical logic and derived disciplines.
@@
Likewise, propositional proof complexity is an area of study that has seen a rapid development over the last two decades, which plays as important a role in the theory of feasible proofs as the role played by the complexity of Boolean circuits in the theory of efficient computations; see, e.g., the celebrated work [Raz15] by Razborov. In most cases, according to [Raz15], the basic question of propositional proof complexity can be reduced to that given a mathematical statement encoded as a propositional tautology $\psi$ and a class of admissible mathematical proofs formalized as a propositional proof system $P$, what is the minimal possible complexity of a $P$-proof of $\psi$? In other words, propositional proof complexity aims to understand and analyze the computational resources required to prove propositional tautologies, in the same way that circuit complexity studies the resources required to compute Boolean functions. In light of the point of view given in [Raz15], the task of proving lower bounds for strong proof systems like Frege or Extended Frege modulo any hardness assumption in the purely computational world may be almost as interesting; see e.g., [Raz15]. It is worth noting that it has been observed by Razborov [Raz15] that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$ implies lower bounds for any propositional proof system. In addition, see, e.g., [Raz03] for interesting directions in propositional proof complexity.
@@
Let us return back to the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem for this moment. As a matter of fact, the so-called $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem is the central problem in proof complexity; see, e.g., the excellent survey [Kra19]. It formalizes the question of whether or not there are efficient ways to prove the negative cases in ours and in many other similar examples. Moreover, the ultimate goal of proof complexity is to show that there is no universal propositional proof system allowing for efficient proofs of all tautologies, which is equivalent to showing that the computational complexity class $\mathcal{NP}$ is not closed under the complement; see e.g., [Kra19].
Particularly, in 1974, Cook and Reckhow showed in [CR74] that there exists a super proof system if and only if $\mathcal{NP}$ is closed under complement; that is, if and only if $\mathcal{NP}={\rm co}\mathcal{NP}$. It is, of course, open whether any superproof system exists. The above result has led to what is sometimes called “Cook’s program” for proving $\mathcal{NP}\neq{\rm co}\mathcal{NP}$: prove superpolynomial lower bounds for proof lengths in stronger and stronger proposition proof systems until they are established for all abstract proof systems. Cook’s program, which also can be seen as a program to prove $\mathcal{P}\neq\mathcal{NP}$, is an attractive and plausible approach; unfortunately, it has turned out to be quite hard to establish superpolynomial lower bounds on common proof systems; see, e.g., the excellent survey on proof complexity [Bus12].
@@ 1.4. Organization
The remainder of this paper is organized as follows: For the convenience of the reader, we will review some notions in the next section closely associated with our discussions and fix some notation we will use in the following context, with some useful technical lemmas also presented in the next section. In Section 3, we prove that all polynomial-time nondeterministic Turing machines are enumerable and all ${\rm co}\mathcal{NP}$-machines are enumerable, thus showing that there exists an enumeration of all ${\rm co}\mathcal{NP}$-machines. Section 4 contains the definition of our universal nondeterministic Turing machine, which accepts a language $L_{s}$ not in ${\rm co}\mathcal{NP}$, showing the desired lower bounds. Section 5 is devoted to showing that the language $L_{s}$ is in fact in $\mathcal{NP}$, proving the upper bounds required, in which we finish the proof of Theorem 1.1 at our own convenience. In Section 6 we will show Theorem 1.3 to break the so-called Relativization Barrier. In Section 7, we show that there are ${\rm co}\mathcal{NP}$-intermediate languages in the complexity class ${\rm co}\mathcal{NP}$. In Section 8, we prove Theorem 1.9, which says that no Frege proof systems are polynomially bounded, thus answering an important open question in the area of proof complexity. Finally, we list some open problems for future research in the last section.
@@ 2. Preliminaries
Let $\mathbb{N}_{1}=\{1,2,3,\cdots\}$, where $\infty\notin\mathbb{N}_{1}$, i.e., the set of all positive integers. We also denote by $\mathbb{N}$ the set of natural numbers, [^3] For the natural number system and $\infty$ (i.e., $\infty$ here), we refer the reader to the interesting textbook [Tao22], in which Tao provides a full and philosophical treatment of them. i.e.,
@@
means that there exists a positive integer $N_{0}$ and a positive constant $M$ such that
@@
The little $o$ notation also indicates the order of growth of some quantity as a function of $n$ or the limiting behavior of a function, but with different meanings. Specifically, that $T(n)$ is little $o$ of $t(n)$, i.e.,
@@
Throughout this paper, the computational modes used are nondeterministic Turing machines (or their variants, such as nondeterministic Turing machines with oracle). We follow the standard definition of a nondeterministic Turing machine given in the standard textbook [AHU74]. Let us first introduce the precise definition of a nondeterministic Turing machine as follows:
@@
Let $M$ be a nondeterministic Turing machine, and let $w$ be an input. Then $M(w)$ represents that $M$ is on input $w$.
@@
Let $M$ be a polynomial-time (say, $t(n)$ time-bounded) nondeterministic Turing machine over the input alphabet $\{0,1\}$. Then, under the “there exists no accepting path” condition, a word $w$ is rejected by $M$ if there exists no computation path $\gamma$ putting the machine into an accepting state on input $w$ (in such a situation, we say $M$ rejects $w$). Clearly, $M$ will define the language $\overline{L}(M)$, which is the complement of $L(M)$ (i.e., $\overline{L}(M)=\{0,1\}^{*}\setminus L(M)$). Specifically, $w\in\overline{L}(M)$ if and only if $M$ rejects $w$ as the input to $M$.
Since ${\rm co}\mathcal{NP}=\{\overline{L}\,:\,L\in\mathcal{NP}\}$, a polynomial-time nondeterministic Turing machine $M$ accepts the language $L$ under the condition of “there exists at least one accepting path,” then the same machine $M$ will simultaneously define the language $\overline{L}$ under the condition of “there exists no accepting path.” Thus, we can conveniently define the notion of a ${\rm co}\mathcal{NP}$ machine as follows:
@@
A ${\rm co}\mathcal{NP}$ machine $M$ is exactly a polynomial-time nondeterministic Turing machine $M$ working under the “there exists no accepting path” condition. We define the language accepted by the ${\rm co}\mathcal{NP}$ machine $M$ to be $\overline{L}(M)$.  [^4] In short, there is an interesting relationship: ${\rm co}\mathcal{NP}$ machine $M$ accepts the language $\overline{L}(M)$, but polynomial-time nondeterministic Turing machine $M$ rejects $\overline{L}(M)$; polynomial-time nondeterministic Turing machine $M$ accepts the language $L(M)$, but ${\rm co}\mathcal{NP}$ machine $M$ rejects $L(M)$. Namely, for any input $w$, $w\in\overline{L}(M)$ if and only if polynomial-time nondeterministic Turing machine $M$ rejects $w$.
Specifically, assume that the $t(n)$ time-bounded ${\rm co}\mathcal{NP}$ machine $M$ (over the input alphabet $\{0,1\}$) accepts the language $\overline{L}(M)$. Then, clearly, $w\notin\overline{L}(M)$ if and only if polynomial-time (i.e., $t(n)$ time-bounded) nondeterministic Turing machine $M$ accepts $w$.